OpenStudy (anonymous):

Question 2 (Multiple Choice Worth 2 points) Which of the following is a zero for the function f(x) = (x + 3)(x – 7)(x + 5)? x = –7 x = –3 x = 3 x = 5 -------------------------------------------------------------------------------- Question 3 (Multiple Choice Worth 2 points) Find the remainder when f(x) = 5x3 + 7x + 5 is divided by x + 2. –49 –7 11 59 -------------------------------------------------------------------------------- Question 4 (Multiple Choice Worth 2 points) Find the remainder when f(x) = 2x3 – 12x2

OpenStudy (anonymous):

For number 2... you have to use the zero factor theorem and set each factor to zero So either x+3=0, x-7=0 or x+5=0 and when you solve you will see that x=-3 is a zero.

OpenStudy (anonymous):

thanks do u know the others

OpenStudy (anonymous):

Nope, sorry :(.